OpenStudy (anonymous):

find the number of divisors of 21600 excliding 1 and number itself

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

21600=2^5*3^3*5^2 so no of divisors=(5+1)(3+1)(2+1)-1 =71